OpenStudy (anonymous):

can some one help me with what is the solution for 2x+x-11+3-7x=15

OpenStudy (anonymous):

Are there any answer choices?

OpenStudy (anonymous):

Or no.

OpenStudy (anonymous):

yes

OpenStudy (zale101):

start by combining like terms

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What are the answer choices? :[

OpenStudy (anonymous):

x = -23/4

OpenStudy (anonymous):

Yeah, but

OpenStudy (zale101):

\((2x+x-11+3-7x)=15\) \((\color{red}{2x+x-7x}-11+3)=15\)

OpenStudy (anonymous):

thks for ur help

OpenStudy (anonymous):

First you simplify both side

OpenStudy (anonymous):

So -4x -8 = 15

OpenStudy (anonymous):

After simplifing you will add 8 to both sides

OpenStudy (anonymous):

-4x - 8 + 8 = 15 + 8 -4x=23

OpenStudy (anonymous):

Then both sides must be divided by -4 so -4x/4 = 23/-4

OpenStudy (anonymous):

Then you should get your answer

OpenStudy (anonymous):

ok thks

OpenStudy (arabpride):

Please close this post if you're done. Thanks! :D